# FPPL: An Efficient and Non-IID Robust Federated Continual Learning Framework

### v1

@@ V-B Performance Comparison
Specifically, in the experiments outlined in Table I and Figure 3(a), non-IID degree was fixed at $\beta=0.5$, and the 200 classes of ImageNet-R and CUB-200 were partitioned into $T=20$, with each task comprising 10 classes. In Table I, the performance metrics of FPPL approach optimal levels. It is notable that FedL2P consistently exhibits lower average forgetting; however, it is crucial to highlight that its accuracy is noticeably lower. This is due to its diminished learning capacity, making it ineffective at acquiring knowledge from later tasks. Consequently, although it demonstrates lower average forgetting for earlier tasks, it also exhibits lower overall global accuracy, contradicting the primary objective of FCL to mitigate catastrophic forgetting while adapting to new tasks. In contrast, FPPL achieves both high accuracy and low average forgetting, effectively fulfilling the purpose of FCL.
@@
Figure 3(a) supports these empirical findings, illustrating that FedSLCA and GLFC both exhibit higher accuracy in the early stages of the 20-task ImageNet-R when there are fewer classes. This is attributed to the overall fine-tuning of the representation layer, enhancing learning ability, especially when dealing with challenging instances of the ImageNet-R dataset. However, this process causes accuracy to gradually decline with an increasing number of learning tasks. In contrast, FPPL tends to maintain a stable state, consistently achieving high accuracy.
@@ V-C Analysis of Efficiency
This section provides a comprehensive efficiency analysis of FPPL, as summarized in Table II, evaluating it across three dimensions: communication cost, storage requirements, and computational burden.
Communication Cost. For FPPL, the transmitted parameters encompass the task-specific prompt, coslinear, classifier, and prototypes. On 20-task ImageNet-R and 20-task CUB-200, with $T=20$, communication cost per round $COMM=253,640$. For GLFC, it not only needs to transmit the entire updated model but also the gradient of the encoding network (LeNet in the original paper [7]), resulting in a total communication cost $COMM=87,571,816$.
Storage Requirement. FPPL is not exempt from extra storage requirements, specifically $ESTG$ on the client. It necessitates storing the frozen task-specific prompts, with a maximum $ESTG=1,536,000$ after $T=20$. In contrast, GLFC maintains an entire old model for knowledge distillation, which for the ViT-B/16 model results in an extra storage $ESTG=85,952,456$.
TABLE II: Efficiency comparison between GLFC and FPPL on 20-task ImageNet-R and 20-task CUB-200.
$RF?$ indicates whether the method is rehearsal-free.
$COMM$ gives the maximum communication cost per round.
$ESTG$ gives the maximum extra stored model parameter
$PARAM$ gives the size of tunable parameters (excluding $\theta$).
Method $RF?$ $COMM$ ($\downarrow$) $ESTG$ ($\downarrow$) $PARAM$ ($\downarrow$)
@@
Computational Burden. During local training on the client side, FPPL tunes only the task-specific prompt, coslinear, and classifier. With a maximum fine-tuning parameter setting of $T=20$, the total number of tunable parameters is $PARAM=92,160$, which is just $0.11\%$ of the parameters in the fully fine-tuned ViT backbone. Conversely, GLFC typically fine-tunes the entire ViT backbone, resulting in $PARAM=85,798,656$. Additionally, compared to rehearsal-based methods like GLFC, FPPL avoids incorporating data from previous tasks during training on the client side, further reducing computational overhead.
@@ V-D Analysis of Non-IID Robustness
Additionally, we further analyzed FPPL’s robustness to non-IID by evaluating FPPL-WOP, which excludes the prototype pool $\bm{C}_{g}$ and relies solely on disposable local prototype sets $\{{\bm{C}}_{k}^{t}\}_{k=1}^{K}$ for classifier debiasing.
@@
This robustness can be attributed to the prototype-based debiasing mechanism, enabling the classifier to fully leverage local data characteristics. As the prototype pool primarily aims to mitigate performance degradation caused by catastrophic forgetting due to missing old task data, FPPL-WOP retains considerable robustness by temporarily utilizing $\{{\bm{C}}_{k}^{t}\}_{k=1}^{K}$ for classifier debiasing, despite displaying a larger performance drop compared to FPPL.
@@ V-E Ablation Study
To further investigate the impact of individual components of FPPL, we conducted an ablation study on the ImageNet-R dataset, setting $T=20$ and $\beta=0.5$. Table III provides insights into the influence of removing core components: unified representation loss, prompt fusion function, and classifier debiasing mechanism.
Specifically, the exclusion of the unified representation loss results in performance degradation, as it hinders the contrastive learning process, which is crucial for ensuring consistent feature representations across diverse clients by leveraging global prototypes. Similarly, the ablation of the fusion function leads to reduced accuracy and increased average forgetting, signifying the failure to effectively exploit the knowledge encapsulated in all prompts.
@@
Remarkably, the removal of the debiasing mechanism incurs a significant decline in performance, emphasizing its pivotal role in rectifying classifier bias. In FCL scenarios characterized by non-IID and absence of old task data, classifiers are susceptible to performance degradation induced by bias. Our debiasing mechanism, grounded in lightweight prototypes, effectively addressed this challenge.
@@
Moreover, we examine the effect of scaling up the number of clients on performance. As shown in Table V-E, there is a marginal decline of performance with the increasing number of clients. Within the same non-IID degree $\beta=0.5$, the expansion of clients exacerbates the data heterogeneity among them, thereby adversely affecting performance. However, our debiasing mechanism ensures robustness, with a maximum decrease in average accuracy of only $1.82\%$. This underscores the robustness of FPPL to non-IID scenarios, an important aspect in real-world applications.
TABLE III: Ablation results on 20-task ImageNet-R. $\bar{A}$ gives the accuracy averaged over tasks, $A_{T}$ gives the accuracy of all tasks after the last incremental stage, $\bar{F}$ gives the average forgetting.
Method $\bar{A}$ ($\uparrow$) $A_{T}$ ($\uparrow$) $\bar{F}$ ($\downarrow$)
FPPL $\bm{74.33\pm 0.37}$ $\bm{70.84\pm 0.15}$ $5.07\pm 0.27$
w/o Unified Eq. (6) $72.96\pm 0.45$ $68.42\pm 0.40$ $5.20\pm 0.16$
w/o Fusion Eq. (2) $72.74\pm 0.13$ $67.03\pm 0.16$ $6.20\pm 0.34$
w/o Debias Eq. (4) $59.00\pm 1.96$ $52.78\pm 0.99$ $5.60\pm 0.49$
TABLE IV: FPPL’s performance under different number of clients on 20-task ImageNet-R. $\bar{A}$ gives the accuracy averaged over tasks, $A_{T}$ gives the accuracy of all tasks after the last incremental stage, $\bar{F}$ gives the average forgetting.
Number of Clients $\bar{A}$ ($\uparrow$) $A_{T}$ ($\uparrow$) $\bar{F}$ ($\downarrow$)
$K=10$ (default) $\bm{74.33\pm 0.37}$ $\bm{70.84\pm 0.15}$
$K=20$ $73.08\pm 0.10$ $69.08\pm 0.42$ $6.57\pm 0.29$
$K=30$ $72.98\pm 0.34$ $68.40\pm 0.40$ $7.74\pm 0.57$
$K=40$ $73.01\pm 0.50$ $68.18\pm 0.22$ $8.70\pm 0.55$
TABLE V: Results on CIFAR-100. Results are included for IID and non-IID setting. $\bar{A}$ gives the average accuracy, $\bar{F}$ gives the average forgetting and $COMM$ gives the communication cost per round. Results of Fed-CPrompt are copied from the original paper [1].

### v3

@@ V-B Performance Comparison
Specifically, in the experiments outlined in Table I and Figure 3(a), non-IID degree was fixed at $\beta=0.5$, and the 200 classes of ImageNet-R and CUB-200 were partitioned into $T=20$, with each task comprising 10 classes. In Table I, the performance metrics of FPPL approach optimal levels. It is notable that FedL2P consistently exhibits lower average forgetting; however, it is crucial to highlight that its accuracy is noticeably lower.
This is due to its diminished learning capacity, making it ineffective at acquiring knowledge from later tasks. Consequently, although it demonstrates lower average forgetting for earlier tasks, it also exhibits lower overall global accuracy, contradicting the primary objective of FCL to mitigate catastrophic forgetting while adapting to new tasks. In contrast, FPPL achieves both high accuracy and low average forgetting, effectively fulfilling the purpose of FCL.
@@
Figure 3(a) supports these empirical findings, illustrating that FedSLCA and GLFC both exhibit higher accuracy in the early stages on 20-task ImageNet-R when there are fewer classes. This is attributed to the overall fine-tuning of the representation layer, enhancing learning ability, especially when dealing with challenging instances of the ImageNet-R dataset. However, this process causes accuracy to gradually decline with an increasing number of learning tasks. In contrast, FPPL tends to maintain a stable state, consistently achieving high accuracy on 20-task ImageNet-R and CUB-200.
@@ V-C Analysis of Efficiency
This section presents a comprehensive efficiency analysis of FPPL, comparing it with the representative full fine-tuning method GLFC, as summarized in Table II. The evaluation covers three dimensions: communication cost, extra storage, and computational burden.
Communication Cost. For FPPL, the transmitted parameters encompass the task-specific prompt, coslinear, classifier, and prototypes. The total transmitted parameters can be formulated as $COMM=D\times(N_{all}+T+\frac{N_{all}}{T}+L_{p}\times M)+N_{all}$, where $M$ represents the number of layers in which prompts are inserted.
For 10-task CIFAR-100, with $N_{all}=100$ and $T=10$, $COMM=169,060$. For 20-task ImageNet-R and CUB-200, with $N_{all}=200$ and $T=20$, $COMM=253,640$. In the case of GLFC, it requires transmitting not only the entire updated model but also the gradient of the encoding network (LeNet in the original paper [7]), resulting in a total communication cost $COMM=87,571,816$.
Extra Storage. FPPL requires extra storage for frozen task-specific prompts on the client side, which can be formulated as $ESTG=D\times L_{p}\times M\times T$.
For 10-task CIFAR-100, with $T=10$, final extra storage requirement $ESTG=768,000$. For 20-task ImageNet-R and CUB-200, $ESTG=1,536,000$. In contrast, GLFC maintains an entire old model for knowledge distillation, which for the ViT-B/16 model results in $ESTG=85,952,456$.
We further analyze the extra server-side storage, where FPPL maintains local prototypes in a prototype pool. Assuming no missing classes across all clients, the total extra server-side storage requirement after $T$ tasks is $D\times K\times N_{\text{all}}$. When $K=10$, for CIFAR-100 with $N_{\text{all}}=100$, the required storage is $768\times 10\times 100=768,000$ (float). As CIFAR-100 images have a size of $32\times 32\times 3=3072$ (int), the storage equals: $768,000$ floats $\times$ $4$ bytes/float $\div$ $3072$ bytes $=1000$ images. Similarly, for ImageNet-R and CUB-200 with $N_{\text{all}}=200$, and resized images have a size of $224\times 224\times 3=150,528$ (int), the extra server-side storage amounts to: $1,536,000$ floats $\times$ $4$ bytes/float $\div$ $150,528$ bytes $\approx 41$ images.
TABLE II: Client-side efficiency comparison between the full fine-tuning GLFC and prompt-based FPPL on 20-task benchmarks. $RF$ indicates whether the method is rehearsal-free.
Method $RF$ $COMM$ ($\downarrow$) $ESTG$ ($\downarrow$) $PARAM$ ($\downarrow$)
@@
Computational Burden. During local training on the client side, FPPL tunes only the task-specific prompt, coslinear, and classifier. The tunable parameters excluding classifier can be formulated as $PARAM=D\times(L_{p}\times M+T)$.
With a maximum setting of $T=20$, the tunable parameters excluding classifier is $PARAM=92,160$, which is just $0.11\%$ of the parameters in the fully fine-tuned ViT backbone. For 10-task CIFAR-100, with $T=10$, tunable parameters excluding classifier is $PARAM=84,480$. Conversely, GLFC typically fine-tunes the entire ViT backbone, resulting in $PARAM=85,798,656$. Additionally, compared to rehearsal-based methods like GLFC, FPPL avoids incorporating data from previous tasks during training on the client side, further reducing computational overhead.
@@ V-D Analysis of Non-IID Robustness
Additionally, we further analyzed FPPL’s robustness to non-IID by evaluating FPPL-WOP, which excludes the prototype pool and relies solely on disposable local prototype set for classifier debiasing.
@@
This robustness can be attributed to the prototype-based debiasing mechanism, which helps the classifier leverage local data characteristics. While FPPL-WOP experiences a larger performance drop compared to FPPL due to the absence of a prototype pool designed to mitigate catastrophic forgetting, it maintains considerable robustness by temporarily using a local prototype set to counteract non-IID issue.
@@ V-E Ablation Study
To further investigate the impact of individual components of FPPL, we conducted an ablation study on the ImageNet-R dataset, setting $T=20$ and $\beta=0.5$. Table III provides insights into the influence of removing three key components: unified representation loss, fusion function, and debiasing.
@@
Remarkably, the removal of the debiasing mechanism incurs a significant decline in performance, emphasizing its pivotal role in rectifying classifier bias. In FCL scenarios characterized by non-IID and absence of data from previous tasks, classifiers are susceptible to performance degradation induced by bias. Our debiasing mechanism, grounded in lightweight prototypes, effectively addressed this challenge.
Additionally, both the unified representation loss and the fusion function contribute to FPPL. The unified representation loss unifies the representations across clients through contrastive learning between extracted features and global prototypes, thereby improving accuracy and reducing forgetting. Similarly, incorporating the fusion function helps to leverage the knowledge encapsulated in all task-specific prompts, which becomes increasingly crucial as the number of tasks grows.
@@
Moreover, we examine the effect of scaling up the number of clients on performance. As shown in Table V-E, there is a marginal decline of performance with the increasing number of clients. Within the same non-IID degree, the expansion of clients exacerbates the data heterogeneity among them, thereby adversely affecting performance. However, our FPPL demonstrates strong robustness, with a maximum decrease in average accuracy of only $1.82\%$.
TABLE III: Ablation results on three key components of our FPPL framework: unified representation loss (U.), fusion function (F.), and debiasing mechanism (D.). The experiments are conducted on 20-task ImageNet-R.
U. F. D. $\bar{A}$ ($\uparrow$) $A_{T}$ ($\uparrow$) $\bar{F}$ ($\downarrow$)
$54.75\pm 2.60$ $46.78\pm 1.78$ $10.18\pm 0.91$
$\bm{\checkmark}$ $55.89\pm 2.66$ $48.78\pm 1.89$ $8.74\pm 0.81$
$\bm{\checkmark}$ $56.30\pm 2.28$ $50.13\pm 1.67$ $7.30\pm 0.88$
$\bm{\checkmark}$ $\bm{\checkmark}$ $59.00\pm 1.96$ $52.78\pm 0.99$ $5.60\pm 0.49$
$\bm{\checkmark}$ $71.82\pm 0.10$ $65.63\pm 0.08$ $6.45\pm 0.55$
$\bm{\checkmark}$ $\bm{\checkmark}$ $72.74\pm 0.13$ $67.03\pm 0.16$ $6.20\pm 0.34$
$\bm{\checkmark}$ $\bm{\checkmark}$ $72.96\pm 0.45$ $68.42\pm 0.40$ $5.20\pm 0.16$
$\bm{\checkmark}$ $\bm{\checkmark}$ $\bm{\checkmark}$ $\bm{74.33\pm 0.37}$ $\bm{70.84\pm 0.15}$ $5.07\pm 0.27$
TABLE IV: Performance of FPPL with varying number of clients. The experiments are conducted on 20-task ImageNet-R.
$K$ $\bar{A}$ ($\uparrow$) $A_{T}$ ($\uparrow$) $\bar{F}$ ($\downarrow$)
$\bm{10(\textbf{default})}$ $\bm{74.33\pm 0.37}$ $\bm{70.84\pm 0.15}$
$20$ $73.08\pm 0.10$ $69.08\pm 0.42$ $6.57\pm 0.29$
$30$ $72.98\pm 0.34$ $68.40\pm 0.40$ $7.74\pm 0.57$
$40$ $73.01\pm 0.50$ $68.18\pm 0.22$ $8.70\pm 0.55$
TABLE V: Performance comparison between Fed-CPrompt and FPPL on 10-task CIFAR-100. Results of Fed-CPrompt are copied from the original paper [1].